# Superpixel Tokenization for Vision Transformers: Preserving Semantic Integrity in Visual Tokens

### v1

@@ Abstract
Transformers, a groundbreaking architecture proposed for Natural Language Processing (NLP), have also achieved remarkable success in Computer Vision. A cornerstone of their success lies in the attention mechanism, which models relationships among tokens. While the tokenization process in NLP inherently ensures that a single token does not contain multiple semantics, the tokenization of Vision Transformer (ViT) utilizes tokens from uniformly partitioned square image patches, which may result in an arbitrary mixing of visual concepts in a token. In this work, we propose to substitute the grid-based tokenization in ViT with superpixel tokenization, which employs superpixels to generate a token that encapsulates a sole visual concept. Unfortunately, the diverse shapes, sizes, and locations of superpixels make integrating superpixels into ViT tokenization rather challenging. Our tokenization pipeline, comprised of pre-aggregate extraction and superpixel-aware aggregation, overcomes the challenges that arise in superpixel tokenization. Extensive experiments demonstrate that our approach, which exhibits strong compatibility with existing frameworks, enhances the accuracy and robustness of ViT on various downstream tasks. The source codes are available at: https://github.com/jangsoohyuk/SuiT
@@ 1 Introduction
Transformers [51], originally proposed for Natural Language Processing (NLP), have demonstrated exceptional performance across diverse tasks such as machine translation [55, 19, 5], natural language understanding [27, 34], and text generation [41, 31]. The transformative power of Transformers was again attested by the release of large language models [7], which continue to revolutionize NLP as we know it. Similarly, the Vision Transformer (ViT) [14], which extended the Transformer architecture to image recognition, changed the landscape of Computer Vision research. The subsequent research efforts to broaden the usage of ViTs to various visual tasks, including detection [45, 8, 65], segmentation [32, 63, 48], generation [15, 38, 60], and multi-modal applications [42], have made the ViT the dominant architecture in Computer Vision.
@@
In NLP, the success of Transformers is commonly attributed to the attention mechanism, which captures contextual dependencies by modeling the relationship among input texts, which are expressed as tokens. Widely adopted tokenization methods in NLP, such as Byte Pair Encoding [46], WordPiece [47], and SentencePiece [29], first split each sentence into words, which are then broken down into more granular tokens. For instance, when tokenizing “Attention is All You Need,” each word is separated into: [“Attention,” “is,” “All,” “You,” “Need”]. By definition, this initial word-level splitting of sentences prevents multiple words with different meanings from being fused into a single token.
@@
In contrast, fixed-size grid patches form the basis of tokenization in ViT and its variants [14, 50, 35]. As visualized in Figure 1, when naïvely sliced square image patches are used as tokens, a single token can simultaneously contain two or more visual concepts, e.g., eye and fur. Consequently, this approach in ViT diverges from the core principle of tokenization, which is to convert plain inputs into smaller yet more structured subunits for further processing.
@@
In this work, we employ the concept of superpixels to design a tokenization process that yields semantically preserving tokens that contain a single, isolated visual concept. A superpixel is a cluster of connected pixels in an image that share similar properties, such as color, texture, or position, and it is often used as a fundamental unit in image processing [2, 11, 33]. Figure 1 illustrates how leveraging superpixels alters tokenization of ViT. Grouping pixels into superpixels divides an image into cohesive regions that capture similar semantics, thereby preventing the mixture of unrelated semantic information in each token.
@@
Despite the advantage of superpixels in tokenization, the different shapes and sizes of each superpixel present significant challenges in utilizing them as ViT tokens [25, 59, 62]. In ViTs, tokens are generated from fixed-size square patches at predefined locations through simple flattening and linear projection. Because superpixels consist of varying numbers of pixels, de-facto flattening and linear projection operations in ViT are incompatible with superpixels. Moreover, the location of each superpixel changes dynamically depending on the image. The variability of superpixel locations makes it infeasible to directly apply the positional embeddings of ViT, which are designed for fixed locations. The inadequacy of original ViT operations necessitates novel techniques for addressing irregular geometry and composition to reap the inherent benefits of superpixels.
@@
We thus propose a novel tokenization pipeline that overcomes the aforementioned problems and enables the effective incorporation of superpixels into ViT tokenization. The proposed pipeline consists of two major technical components: pre-aggregate feature extraction and superpixel-aware feature aggregation. The first stage prepares pre-aggregate features for the subsequent superpixel-aware aggregation, which utilizes pooling operations to remove the irregularity and variability of superpixels. Applying pooling operations directly on the input image may result in a severe loss of information. By training the pre-aggregate feature extractor and applying it prior to our feature aggregation, we effectively sidestep this potential information loss. Collectively, this two-stage process allows the proposed tokenization approach to leverage the benefits of superpixels while preserving important details of an image.
@@
We demonstrate the superiority of our Superpixel-Tokenized Vision Transformer (SuiT) compared to the baselines. SuiT outperforms in a variety of tasks including ImageNet-1K classification [12], segmentation [56], transfer learning [66], and self-supervised learning [9]. We further analyze the properties that emerge using our tokenization pipeline by conducting extra examinations. Furthermore, we visualize that our proposed tokenization method indeed results in image tokens that preserve semantics more effectively than the grid-based tokenization method.
@@
  • We propose an effective superpixel-based tokenization method that overcomes the challenges associated with the use of superpixels.
  • We empirically show that our tokenization method produces tokens that better preserve semantic information than grid-based tokens.
  • We demonstrate the effectiveness of our approach through experiments on image classification, transfer learning, self-supervised learning, and zero-shot segmentation.
@@ 2 Related Work
Superpixels, which are perceptually meaningful clusters of pixels, have been a foundational concept in computer vision [2, 11, 33]. Previous studies have attempted to use superpixels in several tasks such as segmentation [30, 37] and object detection [18], but these works did not demonstrate the broad applicability of superpixels across a range of vision tasks.
Another recent work, STViT [23], applies superpixel-like clustering across all layers via supertoken attention, but operates at the token level, resulting in a coarser granularity than superpixels. CoC [36], borrows the concept of superpixels within neural networks but lacks further exploration of superpixel tokenization. Similarly, SPFormer [37] utilizes superpixel-based representations but requires specialized attention modules making it difficult to integrate with existing ViT architectures in a plug-and-play manner.
In contrast to these approaches, our research explores a simple yet effective superpixel tokenization method that can be easily employed in a transformer architecture for a range of general-purpose tasks.
@@
While most ViT variants focus on improving backbone architectures and attention mechanisms using naive square-shaped patches, there have been several attempts in exploration of advanced tokenization strategies tailored for vision tasks. Quadformer [44] and MSViT [16] both propose adaptive tokenization strategies that dynamically adjust token resolutions based on image content.
The studies SPiT [1] and sViT [28] are the most similar to our work in that they target only the tokenizing module while keeping the backbone intact and share similar research motivations. SPiT [1] focused on analyzing this tokenization from the explainable AI (XAI) perspective but did not explore its generalizability across various tasks. Also, SPiT [1] suffers from performance degradation without gradient features, as they lack of abundant features due to their learning free feature extraction. Moreover, its applicability to other learning paradigms, such as self-supervised learning [9], was not investigated.
sViT [28] shared a similar motivation in aiming for semantically preserving tokenization, but it still divides tokens based on a naïve bounding box, failing to create truly semantically preserving tokens. Additionally, resizing these tokens into square shapes compromised the visual structure of objects within the tokens. These shortcomings in design leads leads to limited performance in various tasks.
As described in the previous section, we propose a semantically preserving tokenization through superpixels in vision and investigate the performance of our method across various tasks and explore its adaptability to self-supervised frameworks.
Figure 2: Overview of our superpixel tokenization pipeline. Local features are extracted and combined with positional encodings, followed by superpixel-aware aggregation using mean and max pooling to produce superpixel tokens, which are fed into a Vision Transformer.
@@ 3 Preliminaries
In this section, we introduce the key concepts and notations that are integral to our work, which aims to construct semantically preserving ViT tokens with superpixels. In Section 3.1, we describe the mechanism of superpixel algorithms, and in Section 3.2, we go over the traditional ViT tokenization. Subsequently, we discuss the research challenges in adopting superpixels as tokens in ViTs.
@@ 3.1 Superpixel Algorithms
Superpixels are clusters of image pixels grouped together based on color, position, or other similarity measures. They provide a compact and meaningful representation of the image by grouping pixels into regions that align with its underlying structure.
@@
where $\textbf{x}^{(h,w)}$ is the pixel at $(h,w)$. Each superpixel $\mathcal{C}_{k}$ consists of varying numbers of pixels with different shapes and sizes.
@@ 3.2 Tokenization in Vision Transformers
In grid-based tokenization of conventional ViT and its variants [50, 35, 42], the input image $\mathbf{x}$ is divided into $N=\frac{H\times W}{P^{2}}$ non-overlapping patches of size $P\times P$, each represented as $\mathbf{x}^{(i)}\in\mathbb{R}^{P\times P\times C}$, where $\mathit{i}\in\{1,...,N\}$ corresponds to the index of each patch. Each patch $\mathbf{x}^{(i)}$ is flattened into a vector $\mathbf{x}_{p}^{(i)}\in\mathbb{R}^{P^{2}C}$ and linearly projected using an embedding matrix $\mathbf{E}\in\mathbb{R}^{P^{2}C\times D}$, where $D$ is the dimension of tokens. In addition to this, positional embedding $\mathbf{E}^{(i)}_{pos}\in\mathbb{R}^{D}$ of its location is injected to form the input token $\mathbf{z}^{(i)}=\mathbf{x}_{p}^{(i)}\mathbf{E}+\mathbf{E}^{(i)}_{pos}\in\mathbb{R}^{D}$. The set of tokens $\mathbf{Z}=\{\mathbf{z}^{(i)}\mid i\in\{1,2,...,N\}\}$ which originates from patches as above, is given to the ViT models along with a learnable classification token $\mathbf{z}_{\text{cls}}$ for classification.
@@
The tokenization process in ViT requires square-shaped and fixed-sized patches, acquired at uniform locations. These constraints on image patches make a naïve application of superpixels to ViT tokenization challenging. First, the straightforward flatten-and-project operation is incompatible with superpixels that come in varying sizes and shapes [25, 59, 62]. Flattening of superpixels with diverse shapes and varying numbers of pixels yields vectors of varying lengths, which hinder a simple linear projection. Second, unlike grid patches with fixed positions, superpixels are located at different positions depending on the image. Consequently, the standard positional embedding, which relies on fixed positions, is inappropriate for superpixels.
@@ 4 Method
To address the aforementioned issues in Section 3, we propose a novel superpixel tokenization pipeline that effectively embeds each superpixel into a single token. The overview of superpixel tokenization is illustrated in Figure 2. Our proposed pipeline incorporates two components: pre-aggregate feature extraction (Section 4.1) and superpixel-aware feature aggregation (Section 4.2). Our aggregation is based on pooling operations to address the inconsistent size and shape of superpixels. However, naïve application of such aggregation on the RGB space may lead to critical loss of information. To sidestep this potential risk, we train a pre-aggregate feature extractor and use its output as input for the superpixel-aware aggregation. These two processes of the proposed tokenization method together exploit the benefits of superpixel-based tokenization while conserving critical image information.
4.1 Pre-aggregate Feature Extraction
Local Feature Extraction
@@ 4.1 Pre-aggregate Feature Extraction
Positional features play a crucial role in transformers, as they are unaware of the order of input tokens. While the original Transformer proposed to use sinusoidal positional encodings with pre-defined frequencies [51], ViTs opted to use learnable positional embeddings for each pre-defined locations [14, 50].
Superpixels have complex positional information of high granularity. Using the learnable positional embeddings as in ViT is inadequate, as it would require excessive number of parameters, i.e., $H\times W$ learnable embeddings each of $D$-dimensions. To keep it parameter-efficient, we adopt the sinusoidal positional encoding with learnable frequencies, which is known to capture high frequency information efficiently [49]. To our knowledge, this positional encoding scheme has not been applied to the Transformer literature before, and is a peculiar choice to adapt to our design of using superpixels of high-granularity.
@@
Given the local features $\mathbf{F}_{local}$ and the positional features $\mathbf{F}_{pos}$, we integrate the two to obtain our pre-aggregate features. The two feature maps $\mathbf{F}_{local}$ and $\mathbf{F}_{pos}$ are concatenated and passed to a linear projection layer $\mathcal{P}$:
$\mathbf{F}=\mathcal{P}(\langle\mathbf{F}_{local},\mathbf{F}_{pos}\rangle),$ (4)
@@
4.2 Superpixel-aware Feature Aggregation
To aggregate the extracted pre-aggregate features within each superpixel into a representative embedding, we employ average and max pooling, which effectively handles features of varying numbers. The pre-computed superpixel index map $\mathcal{I}$ of an input image x is used as the guideline for determining which features belong to which superpixels.
Given the $D/2$-dimensional pre-aggregate features, we perform average pooling and max pooling within each superpixel cluster. This results in two feature vectors per superpixel: average features $\mathbf{z}_{\text{avg}}$ and max features $\mathbf{z}_{\text{max}}$, each of $D/2$-dimensions:

### v3

@@ Abstract
Transformers, a groundbreaking architecture proposed for natural language processing (NLP), have also achieved remarkable success in computer vision. A cornerstone of their success lies in the attention mechanism, which models relationships among tokens. While the tokenization process in NLP inherently ensures that each token maintains semantic integrity without containing multiple meanings, the grid-based tokenization of Vision Transformer (ViT) relies on uniformly partitioned square image patches, which may result in an arbitrary mixing of visual concepts within a token. In this work, we propose a novel tokenization pipeline that replaces the grid-based tokenization with superpixels, encouraging each token to capture a distinct visual concept. Unlike square image patches, superpixels are formed in varying shapes, sizes, and locations, making direct substitution challenging. To address this, our pipeline first generates pixel-level embeddings and efficiently aggregates them within superpixel clusters, producing superpixel tokens that seamlessly replace patch tokens in ViT. Extensive experiments demonstrate that our approach enhances the performance of ViT on various downstream tasks and introduces intriguing properties such as adaptive inference and semantic integrity in tokens.
@@ 1 Introduction
Transformers [58], originally proposed for natural language processing (NLP), have demonstrated exceptional performance across diverse tasks such as machine translation [62, 22, 5], natural language understanding [30, 38], and text generation [46, 34]. The transformative power of Transformers was again attested by the release of large language models [7, 56]. Similarly, the Vision Transformer (ViT) [14], which extended the Transformer architecture to image recognition, changed the landscape of computer vision research. The subsequent research efforts to broaden the usage of ViTs to various visual tasks, including detection [50, 8, 74], segmentation [35, 72, 53], generation [15, 42, 69], and multi-modal applications [47], have made the ViT the dominant architecture in computer vision.
@@
In NLP, the success of Transformers is commonly attributed to the attention mechanism, which captures contextual dependencies by modeling the relationship among input texts, which are expressed as tokens. The widely adopted tokenization methods in NLP, such as Byte Pair Encoding [51], WordPiece [52] and SentencePiece [32], first divide each sentence into words based on whitespaces. For instance, when tokenizing “John loves Mary,” each word is separated into: [“John”, “ ”, “loves,” “ ”, “Mary”]. This initial word-level splitting of sentences prevents multiple words with different meanings from being fused into a single token, preserving semantic integrity in tokens.
@@
In contrast, in computer vision, fixed-size grid patches form the basis of tokenization in ViT and its variants [14, 55, 39]. As visualized in Figure 1, when naïvely sliced square image patches are used as tokens, a single token can simultaneously contain two or more visual concepts, e.g., eye and fur. This is analogous to splitting “John loves Mary” into arbitrary fixed-size fragments like [“Joh,” “n l,” “ove”, “s M”, “ary”], disregarding semantic boundaries. Consequently, this grid-based tokenization deviates from the semantic integrity inherent in NLP tokenization, leading to tokens with arbitrarily mixed visual concepts.
@@
In this work, we employ the concept of superpixels to design a tokenization process. A superpixel [2, 11, 36] is a cluster of connected pixels in an image that share similar properties, such as color, texture, or position, and is often used as a unit of semantic representation in various studies [17, 16]. Building on this property, superpixels emerge as a viable alternative to square tokens for preserving semantic integrity. Figure 1 illustrates how leveraging superpixels alters tokenization of ViT. Grouping pixels into superpixels divides an image into cohesive regions that capture similar semantics, thereby minimizing the mixture of unrelated semantic information in each token.
@@
Despite the potential advantages of superpixels in tokenization, the different shapes and sizes of each superpixel present significant challenges in utilizing them as a direct replacement for visual tokens [28, 67, 71]. In ViTs, tokens are generated from fixed-size square patches at predefined locations through simple flattening and linear projection operations. Because superpixels consist of varying numbers of pixels, de facto flattening and linear projection operations in ViT are incompatible with superpixels. Moreover, the location of each superpixel changes dynamically depending on the image. The variability of superpixel locations makes it infeasible to directly apply the positional embeddings of ViT, which are designed for fixed locations. The inadequacy of original ViT operations necessitates novel techniques to address irregular geometry and composition to reap the inherent benefits of superpixels.
@@
We thus propose a novel tokenization pipeline that overcomes the aforementioned problems and enables the effective incorporation of superpixels into ViT tokenization. The proposed pipeline consists of two major technical components: pixel-level embedding and superpixel-aware aggregation. The first stage generates pixel-level embeddings consisting of local and positional features. Second, these pixel-level embeddings are then aggregated according to superpixel clusters. The aggregation process employs pooling operations to remove the irregularity and variability of superpixels. Collectively, this two-stage process allows the proposed tokenization approach to leverage the benefits of superpixels while preserving important details of an image.
@@
We demonstrate the superiority of our Superpixel-Tokenized Vision Transformer (SuiT) compared to other baselines. SuiT outperforms in a variety of tasks including ImageNet-1K classification [12], transfer learning [75], and zero-shot segmentation [64]. We further analyze intriguing properties that emerge using our tokenization pipeline, such as adaptive inference and semantic integrity.
@@
  • We propose an effective superpixel-based tokenization pipeline that overcomes the challenges associated with the use of superpixels.
  • We empirically show that our tokenization pipeline induces notable properties including adaptive inference and semantic integrity in tokens.
  • We demonstrate the effectiveness of our approach through experiments on image classification, transfer learning, and zero-shot segmentation.
@@ 2 Related Work
Superpixels, which are perceptually meaningful clusters of pixels, have been a foundational concept in computer vision [2, 11, 36]. Previous studies have attempted to use superpixels in several tasks such as segmentation [33, 41] and object detection [21].
A recent work, STViT [26], applies superpixel-like clustering named supertoken attention in the Transformer backbone. However, unlike superpixels which are formed at the pixel level, clusters in STViT are formed at the square token level, resulting in coarser granularity. CoC [40] borrows the concept of superpixels within neural networks, but does not actually exploit superpixels in the network. Similarly, SPFormer [41] utilizes superpixel-based representations but requires specialized attention modules making it difficult to integrate with existing ViT architectures in a plug-and-play manner.
These works all modify the transformer backbone to incorporate the concept of superpixels but do not explore their use in tokenization. In contrast, our work introduces a simple yet effective superpixel tokenization method that can be seamlessly integrated into the Transformer architecture without altering the backbone for a range of general-purpose tasks.
@@
While most ViT variants focus on improving backbone architectures and attention mechanisms using naïve square image patches, there have been several attempts in exploration of advanced tokenization strategies tailored for vision tasks. Quadformer [49] and MSViT [19] both propose adaptive tokenization strategies that dynamically adjust token resolutions based on the image content.
The studies SPiT [1] and sViT [31] are the most similar to our work as they focus solely on the tokenization module while keeping the backbone intact and share similar research motivations. SPiT [1] uses superpixel tokenization but focuses on analyzing the tokenization from the perspective of explainable AI but do not demonstrate favorable performance compared to original ViT.
sViT [31] shares a similar motivation in preserving semantic integrity during tokenization, but it still divides tokens based on a naïve bounding box, failing to create tokens with semantic integrity. Additionally, resizing these tokens into square shapes compromised the visual structure of objects within the tokens. These shortcomings in design leads to limited performance in various tasks.
As described in Section 1, we propose a tokenization that encourages semantic integrity in tokens through superpixels. We investigate the performance of our method across various tasks and explore its intriguing properties.
Figure 2: Overview of our superpixel tokenization pipeline. Local features are extracted and combined with positional encodings, followed by superpixel-aware aggregation using average and max pooling to produce superpixel tokens, which are fed into Vision Transformer.
@@ 3 Preliminaries
In this section, we introduce the key concepts and notations that are integral to our work, which aims to construct ViT tokens that preserve semantic integrity using superpixels. In Section 3.1, we describe the mechanism of superpixel algorithms, and in Section 3.2, we go over the traditional ViT tokenization. Subsequently, we discuss the research challenges in adopting superpixels as tokens in ViTs.
@@ 3.1 Superpixel Algorithms
Superpixels are clusters of image pixels grouped together based on color, position, or other similarity measures. [2, 11, 36] They provide a compact and meaningful representation of the image by grouping pixels into regions that align with its underlying structure.
@@
where $\textbf{x}^{(h,w)}$ is the pixel at $(h,w)$. Each superpixel $\mathcal{C}_{k}$ consists of varying numbers of pixels with irregular spatial distributions.
@@ 3.2 Tokenization in Vision Transformers
In grid-based tokenization of conventional ViT and its variants [55, 39, 47], the input image $\mathbf{x}$ is divided into $N=\frac{H\times W}{P^{2}}$ non-overlapping patches of size $P\times P$, each represented as $\mathbf{p}^{(i)}\in\mathbb{R}^{P\times P\times C}$, where $\mathit{i}\in\{1,...,N\}$ corresponds to the index of each patch. Each patch $\mathbf{p}^{(i)}$ is flattened into a vector $\mathbf{p}_{flat}^{(i)}\in\mathbb{R}^{P^{2}C}$ and linearly projected using an embedding matrix $\mathbf{E}\in\mathbb{R}^{P^{2}C\times D}$, where $D$ is the size of hidden dimension. In addition to this, positional embedding $\mathbf{E}^{(i)}_{pos}\in\mathbb{R}^{D}$ of its location is injected to form the input token $\mathbf{z}^{(i)}=\mathbf{p}_{flat}^{(i)}\mathbf{E}+\mathbf{E}^{(i)}_{pos}\in\mathbb{R}^{D}$. The set of tokens $\mathbf{Z}=\{\mathbf{z}^{(i)}\mid i\in\{1,2,...,N\}\}$ which originates from patches as above, is given to the ViT models along with a learnable classification token $\mathbf{z}_{\text{cls}}$ for classification.
@@
The tokenization process in ViT requires square-shaped and fixed-sized patches, acquired at uniform locations. These constraints on image patches make a naïve application of superpixels to ViT tokenization challenging. First, the straightforward flatten-and-project operation is incompatible with superpixels that come in varying sizes and shapes [28, 67, 71]. Flattening of superpixels with diverse shapes and varying numbers of pixels yields vectors of varying lengths, which hinder a simple linear projection. Second, unlike grid patches with fixed positions, superpixels are located at different positions depending on the images. Consequently, the standard positional embedding, which relies on fixed positions, is inappropriate for superpixels.
@@ 4 Method
To address the issues in Section 3, we propose a superpixel tokenization pipeline that embeds each superpixel into a single token. The overview of superpixel tokenization is illustrated in Figure 2. Our pipeline consists of two components: Pixel-level Embedding (Section 4.1) and Superpixel-aware Aggregation (Section 4.2). Our aggregation is based on pooling operations to address the varying sizes and shapes of superpixels. However, naïve application of such aggregation on the RGB space may lead to critical loss of information. (Table 4) To sidestep this potential risk, we use a pixel-level embedding layer to enrich pixel-level information before superpixel-aware aggregation. These two processes of the proposed tokenization method together exploit the benefits of superpixel-based tokenization while conserving critical image information.
4.1 Pixel-level Embedding
Local Features
@@ 4.1 Pixel-level Embedding
Positional features play a crucial role in Transformers, as they are unaware of the order of input tokens. While the original Transformer proposed to use sinusoidal positional encodings (PE) with pre-defined frequencies [58], ViTs opted to use learnable positional embeddings for each pre-defined locations [14, 55].
Superpixels have complex positional information of high granularity. Using the learnable positional embeddings along every pixels as in ViT is inadequate, as it would require excessive number of parameters, i.e., $H\times W$ learnable embeddings each of $D$-dimensions. To keep it parameter-efficient, we adopt the sinusoidal positional encoding with learnable frequencies, which is known to capture high frequency information efficiently [54]. To our knowledge, this positional encoding scheme has not been applied to the Transformer literature before, and is a peculiar choice to adapt to our design of using superpixels of high-granularity.
@@
Given the local features $\mathbf{F}_{local}$ and the positional features $\mathbf{F}_{pos}$, we integrate the two to obtain our pixel-level embeddings. The two feature maps $\mathbf{F}_{local}$ and $\mathbf{F}_{pos}$ are concatenated and passed through a linear projection layer, represented as a weight matrix $\mathbf{W}_{P}$:
$\mathbf{F}=\langle\mathbf{F}_{local},\mathbf{F}_{pos}\rangle\mathbf{W}_{P}$ (4)
@@
4.2 Superpixel-aware Aggregation
To form an embedding of each superpixel, we aggregate the pixel-level embeddings within each superpixel. The superpixel index map $\mathcal{I}$, acquired by the superpixel algorithm, serves as a guideline in determining which embeddings correspond to which superpixels. Based on $\mathcal{I}$, we aggregate $D/2$-dimensional pixel-level embeddings within each superpixel via average and max pooling, which effectively handles embeddings of varying numbers. This results in two embedding vectors per superpixel: average embeddings $\mathbf{z}_{\text{avg}}$ and max embeddings $\mathbf{z}_{\text{max}}$, each of $D/2$-dimensions: